40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 26 02:15:36 2020, gautam, Update, LSC, Lock acquisition portal entry 
    Reply  Tue May 26 02:31:00 2020, gautam, Update, LSC, Lock acquisition sequence PRFPMIlock_1274418200_1274418550.pdf
       Reply  Wed May 27 20:14:51 2020, Koji, Update, LSC, Lock acquisition sequence 
    Reply  Tue May 26 02:37:19 2020, gautam, Update, LSC, DARM loop measurement and fitting DARM_TF.pdfDARM_TF_breakdown.pdf
    Reply  Tue May 26 03:01:35 2020, gautam, Update, LSC, CARM loop  CARM_OLTF.pdf
       Reply  Wed Jun 3 01:46:14 2020, gautam, Update, LSC, CARM loop  CM_loop_topology.pdfCARM_TFs.pdfCARM_OLTF.pdfCARM_xover.pdfCARM_OLG_evolution.pdf
    Reply  Tue May 26 03:06:59 2020, gautam, Update, LSC, PRFPMI sensing matrix PRFPMI_20200524sensMat.pdfPRFPMI_20200524sensMatHistograms.pdf
    Reply  Tue May 26 03:26:58 2020, gautam, Update, LSC, Preliminary noise budget PRFPMI_NB.pdf
    Reply  Tue May 26 14:32:44 2020, gautam, Update, LSC, Arm transmission RIN armRIN.pdf
       Reply  Wed May 27 19:36:33 2020, Koji, Update, LSC, Arm transmission RIN 
          Reply  Thu May 28 18:36:45 2020, gautam, Update, LSC, Arm transmission RIN PRFPMIcorner_ASC_PIT_1274419354_1274419654.pdfPRFPMIcorner_ASC_YAW_1274419354_1274419654.pdfPRFPMIcorner_ASC_coherence_1274419354_1274419654.pdf
             Reply  Fri May 29 00:34:57 2020, rana, Update, LSC, Arm transmission RIN 
             Reply  Wed Jun 3 02:14:32 2020, gautam, Update, ASC, PRC ASC improves arm transmission RIN PRC_ASCsignals.pdfarmRIN_PRC_ASC.pdfPRFPMIcorner_ASC_PIT_1275190251_1275190551.pdfPRFPMIcorner_ASC_YAW_1275190251_1275190551.pdfPRFPMIcorner_ASC_coherence_1275190251_1275190551.pdf
                Reply  Fri Jun 19 16:30:09 2020, gautam, Update, ASC, Some thoughts about ASC sensingResponse.pdfsensingResponse_torque.pdf
                   Reply  Tue Jul 7 14:06:10 2020, gautam, Update, ASC, Some more thoughts about ASC ITM_OL_DCcoupling.png
    Reply  Tue May 26 16:00:06 2020, gautam, Update, LSC, Power buildup diagnostics PRFPMIcorner_DC_1274419354_1274419654.pdfPRFPMIcorner_SB_1274419354_1274419654.pdf
       Reply  Wed May 27 17:41:57 2020, Koji, Update, LSC, Power buildup diagnostics 
          Reply  Wed Jun 3 02:08:00 2020, gautam, Update, LSC, Power buildup diagnostics PRFPMIcorner_DC_1275190251_1275190551.pdfPRFPMIcorner_SB_1275190251_1275190551.pdf
    Reply  Wed Jun 3 01:34:53 2020, gautam, Update, LSC, Lock acquisition update portal 
       Reply  Wed Jun 3 03:29:26 2020, Koji, Update, LSC, Lock acquisition update portal 
          Reply  Wed Jun 3 11:40:56 2020, gautam, Update, LSC, Lock acquisition update portal 
             Reply  Wed Jun 3 18:49:47 2020, gautam, Update, LSC, PRG and CARM signal sign armCavReflectivities.pdfIFOreflectivities.pdfPDHerrSigs.pdfPRGvsLoss_finesse.pdf
Message ID: 15352     Entry time: Tue May 26 03:06:59 2020     In reply to: 15348
Author: gautam 
Type: Update 
Category: LSC 
Subject: PRFPMI sensing matrix 

Summary:

The response of the PRFPMI length degrees of freedom as measured in the LSC PDs was characterized. Two visualizations are in Attachment #1 and Attachment #2.

Details:

  • The sensing matrix infrastructure in the c1cal model was used.
  • The oscillator frequencies are set between 300 - 315 Hz.
  • Notch filters at these frequencies were enabled in the CDS filter banks, to prevent actuation at these frequencies (except for CARM, in which case the loop gain is still non-negligible at ~300 Hz, this correction has not yet been applied).
  • Mainly, I wanted to know what the DARM sensing response in AS55_Q is. 
    • The measurement yields 2.3e13 cts/m. This is a number that will be used in the noise budget to convert the measured DARM spectrum to units of m/rtHz.
    • We have to multiply this by 10/2^15 V/ct, undo the 6dB whitening gain on the AS55_Q channel, and undo the ~5x gain from V_RF to V_IF (see Attachment #4 of this), to get ~0.69 GV/m from the RFPD.
    • The RF transimpedance of AS55_Q is ~550 ohms, and accounting for the InGaAs responsivity, I get an optical gain of 1.8 MW/m. Need to check how this lines up with expectations from the light levels, but seems reasonable.
    • Note that T_SRM is 10%, we dump 70% of the output field into the unused OMC, and there is a 50/50 BS splitting the light between AS55 and AS110 PDs. Assuming 90% throughput from the rest of the chain, we are only sensing ~1.3 % of the output DARM field.
  • Apart from this, I can also infer what the matrix elements / gains need to be for transitioning the PRMI control from 3f to 1f signals. To be done...
  • I found these histograms in Attachment #2 to be a cute way of (i) visualizing the variance in the magnitude of the sensing element and (ii) visualizing the separation between the quadratures, which tells us if the (digital) demod phase needs to be modified.
    • The sensing lines were on for 5 minutes (=300 seconds) and the FFT segment length is 5 seconds, so these histograms are binning the 60 different values obtained for the value of the sensing element.
    • The black dashed lines are "kernel density estimates" of the underlying PDFs
    • I haven't done any rigorous statistical analysis on the appropriateness of using this technique for error estimation, so for now, they are just lines...
Attachment 1: PRFPMI_20200524sensMat.pdf  55 kB  Uploaded Tue May 26 04:07:11 2020  | Hide | Hide all
PRFPMI_20200524sensMat.pdf
Attachment 2: PRFPMI_20200524sensMatHistograms.pdf  29 kB  Uploaded Tue May 26 04:07:19 2020  | Hide | Hide all
PRFPMI_20200524sensMatHistograms.pdf
ELOG V3.1.3-